Hi i would like to know the general formula for the sequence
1, 4, 2, 8, 4, 16, 8, 32. I am certain thatit follows a
÷2 X4 sequence.. But please help me for the general formula to answer any nth term. Thanks!

Answer by Alex.33(110)

a{n}=2^(2n-1),n is odd
a{n}=2^(n/2+1), n is even.
It can't be simplified any further. All I did was just treat the odd and even terms independently like two seperate geometric sequences but keep their cardinal n in the original state.
Update: Yes the formula for odd numbers should be a{n}=2^((n-1)/2). There was a typo when I was trying to type in the stuff I wrote on my draft papers.
Sincere apology here.